PH calculation

1.What is the pH at the equivalence point in the titration of a 15.8 mL sample of a 0.483 M aqueous hydrocyanic acid solution with a 0.372 M aqueous barium hydroxide solution? pH =
2.When a 23.4 mL sample of a 0.425 M aqueous hypochlorous acid solution is titrated with a 0.442 M aqueous potassium hydroxide solution, what is the pH at the midpoint in the titration? pH = |
Answer #1

1)

2 HCN + Ba(OH )2----------> Ba(CN)2 + 2 H2O

initially millimoles of HCN = 15.8 x 0.483 = 7.63

3.815 millimoles of Ba(OH )2 must be added to reach equivalence point .

3.815 = V x 0.372

V =10.25 mL Ba(OH )2 must be added

total volume = 15.8 + 10.25 = 26.05 mL

[Ba(CN)2] = 7.63 / 26.05 = 0.292 M

Ba(CN)2 is salt of weak acid for such salts

pH = 1/2 [pKw + pKa + log C]

pKa of HCN = 9.2 standard value

pH = 1/2 [14 + 9.2 + log 0.292]

pH = 1/2 [22.67]

pH = 11.33

pH = 11.3

2.

The pH at the midpoint of the titration (the half-equivalence point) will be equal to the pKa of Hypochlorous acid ( at mid point number of moles of acid is equal to number of moles of its conjugate base)

so the pKa of HOCl = 7.53

so pH = pKa= 7.5


answered by: Ravi Prabhat